OpenStudy (anonymous):

Use a definite integral to find an expression that represents the area of the region between the given curve and the x-axis on the interval [0,b] y=6x^2 answer choices -2b^3, -12b, 12b, 2b^3

OpenStudy (anonymous):

\[\int\limits\limits_{0}^{b}6x^2dx=\left|2x^3 \right|^b_0=2b^3-0=2b^3\]The last one